OpenStudy (anonymous):

Finding the y-intercept of 7x-y = -14

OpenStudy (ashleyisakitty):

First you need to put this into y=ax+b format. Do you know how to do that?

OpenStudy (anonymous):

Yes, I do. I've gotten to the point of where -y = -14 + -7x. The x equals zero and that's where I've gotten stuck.

OpenStudy (ashleyisakitty):

I think the problem is that your plus sign is throwing you off :) When i rearrange it i get -y=-7-14. Now, we are going to have to isolate the y. Do this by dividing each side by -1. -y/-1 = -7-14/-1 What do you get after this?

OpenStudy (ashleyisakitty):

Excuse me, it would actually be -y/-1 = -7x-14/1

OpenStudy (anonymous):

7 x 0 is 0. That's what threw me off because there was a negative sign in front of the 0. That's definitely what confused me. (:

OpenStudy (anonymous):

to find the y intercept, substitute 0 in for x

OpenStudy (ashleyisakitty):

Yes, the next step is to substitute the 0 for x. *cough cough*.... y=7(0)+14 =0+14 = ? and that is your answer :)

OpenStudy (ashleyisakitty):

Excuse me, it would actually be -y/-1 = -7x-14/1

OpenStudy (anonymous):

The 14 was negative, though. I probably sound foolish.

OpenStudy (anonymous):

just sub 0 in for x and you get y = 14

OpenStudy (ashleyisakitty):

No worries Alivia, we all make mistake sometimes!

OpenStudy (mathstudent55):

The equation y = mx + b is the slope-intercept form of the equation of a line. If you write the equation of a line in that form, you can readily find the slope and the y-intercept without any further calculations. The slope is m and the y-intercept is b. If you are asked for the y-intercept of an equation, solve the equation for y, and put in the slope intercept form. Then you can tell quickly what the y-intercept is. In this problem, you were asked to find the y-intercept of 7x - y = -14 You need to solve the equation for y: First, subtract 7x from both sides, and place the -7x to the left of the -14: 7x - y = -14 -7x -7x -------------------- -y = -7x - 14 Since you now have -y, but you want y, you can multiply both sides by -1. -1(-y) = (-1)(-7x) - (-1)(14) y = 7x + 14 From the slope-intercept form you can immediately conclude that the y-intercept is 14. (You can also conclude that the slope is 7, but you were not asked that.)

OpenStudy (anonymous):

Thank you guys so, so much! You've been an amazing help! And oh my goodness, that's very detailed, but it helps a lot. Thank you! <3

OpenStudy (mathstudent55):

You're welcome.